OpenStudy (anonymous):

solve 1/4^(x+1) = 8x?

OpenStudy (mathstudent55):

Is this it? \( \left( \dfrac{1}{4} \right) ^{x + 1} = 8x\)

OpenStudy (anonymous):

yesss

OpenStudy (anonymous):

I don't understand...:(

OpenStudy (mathstudent55):

Is the right side correct. It's 8x, not 8^x?

OpenStudy (anonymous):

whoops its 8^x, sorry

OpenStudy (mathstudent55):

That's better.

OpenStudy (mathstudent55):

\( \left( \dfrac{1}{4} \right)^{x + 1} = 8^x\) Rewrite the left side as a power of 2. Rewrite the right side as a power of 2.

OpenStudy (anonymous):

how do I do that?

OpenStudy (mathstudent55):

\( \left( \dfrac{1}{4} \right) ^ {x + 1} = 8^x\) Start like this. Now multiply the exponents together on each side. \( \left( 2^{-2} \right) ^ {x + 1} = (2^3)^x\)

OpenStudy (anonymous):

x=-2/5?

OpenStudy (mathstudent55):

\(\left( 2^{-2} \right) ^ {x + 1} = (2^3)^x\) \( 2^{-2x -2} = 2^{3x} \) \( -2x -2 = 3x \) \( -5x = 2 \) \(x = -\dfrac{2}{5} \) You are correct.

OpenStudy (anonymous):

yes! thanks :D

OpenStudy (mathstudent55):

wlcm